# Artificial Intelligence HCI Human-Computer Interaction

### v2

@@ 1. Introduction
Recently, Artificial Intelligence (AI) tools have grown increasingly popular, driven by the rapid rise of Large Language Models (LLMs). These powerful models can generate coherent, contextually relevant text across diverse topics and perform tasks such as classification and summarization (Wei et al., 2022a; Bubeck et al., 2023). Large Language Models have been widely adopted across domains, including workspaces, from programming assistance to educational tutoring, and are increasingly finding their way into research contexts.
@@
One area where these models hold significant potential is Qualitative Data Analysis (QDA). Human-Computer Interaction (HCI) researchers are beginning to explore the usage of LLMs throughout their research processes for tasks such as conducting QDA (De Paoli, 2024), generating synthetic data and substituting human participants (Hämäläinen et al., 2023; Schmidt et al., 2024; Argyle et al., 2023), creative ideation (Shaer et al., 2024; Oppenlaender and Hämäläinen, 2023), literature reviews, and writing (Kapania et al., 2024a). Commercial tools have also started integrating AI features (e. g., ATLAS.ti, [^1] https://atlasti.com/ MAXQDA, [^2] https://www.maxqda.com/products/ai-assist and nVivo’s Lumivero AI Assistant [^3] https://lumivero.com/product/nvivo-ai-assistant/(Morgan, 2023).
@@
The rapid adoption of AI tools begs the question: which parts of the QDA workflow are suitable for AI support, and how do we make these determinations? These questions are particularly pivotal for CSCW, where QDA plays a dual role: as a methodological tool, and as a site of collaboration, negotiation, and interpretive practice. Although AI tools can be useful, they raise dilemmas about researcher autonomy (Feuston and Brubaker, 2021) and the risks of imperfect AI outputs (e. g., LLMs hallucinating wrong information) (Xiao and Wang, 2021). As a community, it is imperative that we critically assess which tasks in the QDA workflow are appropriate or inappropriate for AI integration, and to ensure its responsible use. To better understand how AI tools can, and should, be integrated into researchers’ QDA workflows, we conducted semi-structured interviews with HCI researchers about their QDA workflows and opinions on AI integration. We found that while researchers’ QDA workflows share structural similarities, they differ significantly based on project specifics, highlighting the need for AI tools that are highly customizable. Researchers were generally open to AI assistance when it enhanced their workflow but stressed the importance of ethical considerations, trust in AI outputs, and maintaining control.
@@
Our contributions are as follows. First, we provide an empirical account of researchers’ QDA practices, pain points, and challenges. Second, we present a framework for integrating AI into QDA workflows, grounded in researchers’ perspectives on the individual stages of QDA. Based on our findings we ideate tangible scenarios for the integaration of AI, including AI as a productivity tool (e. g., calculating inter-rater reliability), as a mediator (e. g., facilitating consensus), or with human approval (e. g., suggesting codebook components with validation by a researcher). Finally, we discuss how this framework can guide the responsible deployment of AI tools that truly support researchers’ needs and preserve the social value of collaborative analysis practices.
@@ 2.1. Qualitative Data Analysis in CSCW and HCI
Beyond methods, QDA in CSCW and HCI is shaped by interdisciplinary diversity. Researchers bring different epistemological traditions and backgrounds (Fiesler et al., 2019; McDonald et al., 2019), which leads to ongoing debates about transparency, rigor, reliability, and measures thereof, such as Inter-Rater Reliability (IRR),  (Fiesler et al., 2019; McDonald et al., 2019). At the same time, most CSCW and HCI projects involve more than one researcher (McDonald et al., 2019), making QDA a fundamentally collaborative and social practice of negotiating, resolving disagreements, and making sense of ambiguity. This collaboration is often asynchronous and distributed, and supported by digital tools and software (Wallace et al., 2017), increasingly incorporating Natural Language Processing (NLP) and LLM methods. Yet, despite QDA’s centrality, there has been no systematic documentation of how CSCW and HCI researchers actually conduct QDA in practice, or how emerging technologies might reshape it. Our work addresses this gap by synthesizing and extending fragmented conversations into a coherent account of real-world QDA workflows.
@@ 2.2. From Computational Support to AI-Assisted Analysis
While CSCW and HCI researchers draw on a variety of tools for analysis (Fiesler et al., 2019) computer-assisted approaches have long supported QDA. Early tools applied basic NLP techniques such as word frequency analysis, keyword matching, and pattern-based rules (Stone et al., 1962; Marathe and Toyama, 2018). While these methods introduced automation, they lacked the capacity to capture complex human interpretations. Subsequent work introduced supervised machine learning for text classification, extending manual coding into larger datasets (Rietz and Maedche, 2021). Unsupervised approaches such as topic modeling uncovered latent themes without predefined categories (Lennon et al., 2021), and some studies reported surprising alignment with qualitative methods like grounded theory (Baumer et al., 2017; Muller et al., 2016). Such combinations of QDA and machine learning can be especially useful for mixed-methods studies and for analyzing large-scale data, though they raise concerns about transparency and interpretability – values central to qualitative research. Recent advances in generative AI, particularly LLMs like ChatGPT, offer new opportunities. These models are trained on massive text corpora and can generate, summarize, and classify language with a high degree of contextual awareness (Wei et al., 2022b). They encode vast amounts of diverse knowledge, making them potentially adaptable to various tasks within qualitative research. LLMs have shown remarkable performance on annotation tasks where little or no labeled data is provided (Gilardi et al., 2023; Törnberg, 2023a) and have been proposed as collaborators in human annotation teams (Ziems et al., 2024a). Recent prototypes integrate LLMs at different points across the QDA workflow, from inductive coding to codebook development (Bano et al., 2024; Tai et al., 2024; Xiao et al., 2023; Morgan, 2023; Rasheed et al., 2024; Gauthier and Wallace, 2022). This evolution highlights not only the opportunities for scaling qualitative work but also the risks it creates for preserving interpretive rigor.
@@ 2.3. Potentials and Risks of AI in the QDA Workflow
Recent studies show how AI can assist at multiple points in the QDA process. Cuevas et al. (2025) examine the usage of LLMs for conducting user studies and analyzing the collected data. For labeling and annotating, LLM-based systems apply existing codes from a given codebook with notable consistency, though they often miss subtle contextual judgments (Xiao et al., 2023). Lam et al. (2024) develop a LLM-based text analysis tool that derives interpretable concepts from unstructured text, inspired by traditional topic models. For data exploration and inductive coding, LLMs can surface novel patterns and assist in the initial exploration of themes and codes, potentially enabling a deeper understanding. However, while LLMs can replicate many primary themes identified by humans, they often miss more subtle and complex interpretations (Morgan, 2023; Zhang et al., 2023; Torii et al., 2024). Hamilton et al. (2023) observed complementary patterns between human and ChatGPT-generated codes, with each identifying themes that the other did not. Barany et al. (2024) investigated the use of ChatGPT for codebook creation, finding hybrid approaches to be more reliable. They suggest LLMs can contribute to the improvement of codebooks, but stress that human involvement remains crucial. While combining interpretive and positivist methods can cause epistemic tensions and discussions among researchers (Baumer et al., 2017; Muller et al., 2016; Paulus and Marone, 2024), AI also holds potential for mediating discussions and facilitating collaboration by highlighting coding discrepancies, and suggesting resolutions (Gao et al., 2023; Ganji et al., 2018; Drouhard et al., 2017).
@@
While these efforts illustrate potential, they also expose tensions. Concerns persist around bias (Bail, 2024; Törnberg, 2023b; Ziems et al., 2024b; Anis and French, 2023; Owoahene Acheampong and Nyaaba, 2024; Bano et al., 2023), lack of reliability for non-English languages (Heseltine and Clemm von Hohenberg, 2024; Suter and Meckel, 2024), and the risk of treating AI as “culturally agnostic” despite its privileging of certain perspectives (Paulus and Marone, 2024). Such issues are particularly important when engaging Global South contexts, where AI risks reproducing inequalities in representation (Kumar and Karusala, 2021; Wacharamanotham et al., 2020; Rifat et al., 2023; Jang et al., 2023; Okolo et al., 2022). Thus, while AI holds potential to bridge “big” and “small” data approaches and to enhance efficiency, it raises equally urgent questions about reliability, ethics, and researcher agency.
2.4. Designing for Human-AI Collaboration
@@
Integrating AI into QDA is not simply a matter of automation. It requires careful design for human–AI collaboration in interpretive, context-sensitive work (Jiang et al., 2021). Existing tools vary widely in how they support AI involvement, from minimal assistance to active roles in coding and analysis, with different levels of human involvement. Moreover, designing for researchers differing in proficiency in tool usage as well as AI capability (Lubars and Tan, 2019; Feuston and Brubaker, 2021) poses specific challenges, especially in diverse fields such as CSCW and HCI. For example, Gauthier and Wallace (2022) developed a toolkit to facilitate computational thematic analysis, emphasizing the need to design it accessibly for a community with different proficiencies in tool usage as well as AI. Thus, they call for ethical designs to involve friction to remind researchers of the imperfection of computational models and issues concerning ethics and transparency. These challenges are also identified by Kapania et al. (2024b) who found HCI researchers struggling to identify such issues in their projects and needing guidance. Similarly, Lam et al. (2024) suggest tools to be transparent to the researcher and acknowledging their limitations, e. g., by showcasing explicit inclusion criteria and scores. At the same time, designs should avoid creating excessive “unanticipated labor” for researchers (Feuston and Brubaker, 2021).
@@
While prior studies provide valuable insights into the potentials of AI in QDA (Tai et al., 2024; Morgan, 2023; Rasheed et al., 2024; Lam et al., 2024; Gauthier and Wallace, 2022), there is still limited understanding of how AI can be responsibly integrated into researchers’ real-life QDA workflows, respecting their concerns and preferences (Lubars and Tan, 2019; Feuston and Brubaker, 2021). We are thus in need for comprehensive frameworks addressing these issues (Owoahene Acheampong and Nyaaba, 2024; Kapania et al., 2024b). To address these gaps, we study HCI researchers’ QDA workflows and their perspectives on the involvement of AI in qualitative research. Specifically, we investigate the following research questions:
@@ 3.2. Recruitment & Participants
Table 1 summarizes the participant sample. We interviewed 15 HCI researchers with backgrounds in computer science, cybersecurity, engineering, and social sciences, all firmly embedded in HCI research, with some specifying to focus on USP. Twelve participants were early career researchers (mostly PhD candidates), and three were senior researchers. Eight had worked on 1-3 QDA projects, four on 4-6, and three on more than 7 projects, with qualitative experience ranging from 1 to 9 years. We recruited participants until data saturation was reached. We reached data saturation around interview 12 and conducted three additional interviews to confirm this.
@@ 4.1.2. Uncovering Participants’ Actual Coding Workflow
Figure 1. Detailed overview of participants’ real-world coding workflow, consisting of three stages for codebook creation, codebook refinement, codebook application, and different steps taken in inductive and deductive approaches.:

### v3

# arXiv:2501.19275v3[cs.CY] 27 Feb 2026 AI Artificial Intelligence

@@ 1. Introduction
Artificial Intelligence (AI) tools have grown increasingly popular, driven by the rapid rise of Large Language Models (LLMs). These powerful models can generate coherent, contextually relevant text across diverse topics and perform tasks such as classification and summarization (Wei et al., 2022; Bubeck et al., 2023). Large Language Models have been widely adopted across domains, from programming assistance to educational tutoring, and are increasingly finding their way into research contexts.
@@
One area where these models hold significant, yet debated potential is Qualitative Data Analysis (QDA). Human-Computer Interaction (HCI) and CSCW researchers are beginning to explore the usage of LLMs throughout their research processes for tasks such as conducting QDA (De Paoli, 2024), generating synthetic data and substituting human participants (Hämäläinen et al., 2023; Schmidt et al., 2024; Argyle et al., 2023), creative ideation (Shaer et al., 2024; Oppenlaender and Hämäläinen, 2023), literature reviews, and writing (Kapania et al., 2025). Commercial tools have also started integrating AI features (e. g., ATLAS.ti, [^1] https://atlasti.com/ MAXQDA, [^2] https://www.maxqda.com/products/ai-assist and nVivo’s Lumivero AI Assistant [^3] https://lumivero.com/product/nvivo-ai-assistant/(Morgan, 2023).
@@
The rapid adoption of AI tools begs the question: which steps of the QDA workflow are suitable for AI support, and how do we make these determinations? These questions are particularly pivotal for CSCW, where QDA plays a dual role: as a methodological tool, and as a site of collaboration, negotiation, and interpretive practice in interdisciplinary teams with different epistemic traditions (Bjørn and Boulus-Rødje, 2015; Fiesler et al., 2019; McDonald et al., 2019). Although AI tools can be useful, they raise dilemmas about researcher autonomy (Feuston and Brubaker, 2021) and the risks of imperfect AI outputs (e. g., LLMs hallucinating wrong information) (Xiao and Wang, 2021). As a community, it is imperative to critically assess which tasks in the QDA workflow are appropriate or inappropriate for AI integration, and to ensure its responsible use. To better understand how AI tools can, and should (not), be integrated into QDA, we conducted semi-structured interviews with HCI researchers about their QDA workflows and opinions on AI integration. We found that while researchers’ QDA workflows share structural similarities, they differ significantly based on project specifics, highlighting the need for AI tools that are highly customizable. Researchers were generally open to AI assistance when it enhanced their workflow but stressed the importance of ethical considerations, trust in AI outputs, and maintaining control.
@@
Our contributions are as follows. First, we provide an empirical account of researchers’ QDA practices, pain points, and challenges. Second, we present a framework for integrating AI into QDA workflows, grounded in researchers’ perspectives on the individual stages of QDA and prior work, extending beyond the task of coding itself. Based on our findings we ideate tangible scenarios for the integration of AI, including AI as a productivity tool (e. g., taking on more technical tasks), as a mediator (e. g., facilitating consensus), or with human approval (e. g., suggesting codebook components with validation by a researcher). Finally, we discuss how this framework can guide the responsible deployment of AI tools that truly support researchers’ needs and preserve the social value of collaborative analysis practices.
@@ 2.1. Qualitative Data Analysis in CSCW and HCI
Underlying these methods are different epistemologies, from more positivist ones aiming at recognition and discovery, to more interpretive ones, requiring in-depth analysis and centering researcher subjectivity and reflexivity (Arthur and Clark, 2023; Greckhamer and Cilesiz, 2022; Chatzichristos, 2025; Ashcroft, 2023; Soden et al., 2024). Within CSCW and HCI, projects commonly involve more than one researcher (McDonald et al., 2019; Jiang et al., 2021), making QDA a fundamentally collaborative and social practice of negotiating, resolving disagreements, and making sense of ambiguity. This collaboration is often asynchronous and distributed, and supported by digital tools and software (Wallace et al., 2017), increasingly incorporating Natural Language Processing (NLP) and LLM methods into the QDA process – thus, QDA fundamentally becomes computer-supported cooperative work. At the same time, QDA in CSCW and HCI is shaped by interdisciplinary diversity (Bjørn and Boulus-Rødje, 2015). Researchers bring different epistemological traditions and backgrounds (Fiesler et al., 2019; McDonald et al., 2019; Bjørn and Boulus-Rødje, 2015), which leads to ongoing debates about transparency, rigor, reliability, and measures thereof, such as Inter-Rater Reliability (IRR),  (Fiesler et al., 2019; McDonald et al., 2019; Bjørn and Boulus-Rødje, 2015; Soden et al., 2024). Yet, despite QDA’s centrality to CSCW and HCI, there has been no systematic documentation of how CSCW and HCI researchers actually conduct QDA in practice, or how emerging technologies might reshape it. The closest analysis we could find is the work by McDonald et al.(McDonald et al., 2019) documenting how interrater reliability is used in CSCW and CHI. They find that reporting IRR is rare and authors often use other methods to report on reliability. A more comprehensive overview of the whole QDA process, though, is still lacking. Our work addresses this gap by synthesizing and extending fragmented conversations into a coherent account of real-world QDA workflows.
@@ 2.2. From Computational Support to AI-Assisted Analysis
While nowadays, CSCW and HCI researchers draw on a variety of tools for analysis (Fiesler et al., 2019), computer-assisted approaches have long supported QDA, from supporting basic coding tasks, to experiments with automation: Early tools applied basic NLP techniques such as word frequency analysis, keyword matching, and pattern-based rules (Stone et al., 1962; Marathe and Toyama, 2018). While these methods introduced automation, they lacked the capacity to capture complex human interpretations. Subsequent work introduced supervised machine learning for text classification, extending manual coding into larger datasets (Rietz and Maedche, 2021). Unsupervised approaches such as topic modeling uncovered latent themes without predefined categories (Lennon et al., 2021), and some studies reported surprising alignment with qualitative methods like grounded theory (Baumer et al., 2017; Muller et al., 2016). Such combinations of QDA and machine learning can be especially useful for mixed-methods studies and for analyzing large-scale data, though they raise concerns about transparency and interpretability – values central to qualitative research. Recent advances in generative AI, particularly LLMs like ChatGPT, offer new opportunities. These models are trained on massive text corpora and can generate, summarize, and classify language with a high degree of contextual awareness (Wei et al., 2022). They encode vast amounts of diverse knowledge, making them potentially adaptable to various tasks within qualitative research. Recent prototypes integrate LLMs at different points across the QDA workflow, from inductive coding to codebook development (Bano et al., 2023; Tai et al., 2024; Xiao et al., 2023; Morgan, 2023; Rasheed et al., 2024; Gauthier and Wallace, 2022). LLMs have shown remarkable performance on annotation tasks where little or no labeled data is provided (Gilardi et al., 2023; Törnberg, 2023a) and have been proposed as collaborators in human annotation teams (Ziems et al., 2024), indicating a shift from assistance to autonomy of computational methods supporting the cooperative work of QDA.
This evolution highlights not only the opportunities for scaling qualitative work, but also the risks it creates for preserving interpretive rigor and methodological integrity. These concerns are shared by HCI and CSCW researchers (Feuston and Brubaker, 2021; Jiang et al., 2021; Lubars and Tan, 2019), showing reservation and opposition to the integration of AI into QDA, or specifying specific use contexts that retain human agency. While these studies were conducted prior to the advent of LLMs, recent work identified a ”generational gap”, with early career researchers being more open, and more experienced researchers more reserved towards the integration of AI into QDA, calling into question AI and QDA literacy as well as structural inequalities that may facilitate attitudes (Chatzichristos, 2025). This necessitates further engagement with shifting researcher attitudes towards AI-supported QDA.
@@ 2.3. Potentials and Risks of AI in the QDA Workflow
Recent studies show how AI can assist at multiple points in the QDA process. Cuevas et al. (2025) examine the usage of LLMs for conducting user studies and analyzing the collected data. For labeling and annotating, LLM-based systems apply existing codes from a given codebook with notable consistency, though they often miss subtle contextual judgments (Xiao et al., 2023). Lam et al. (2024) develop a LLM-based text analysis tool that derives interpretable concepts from unstructured text, inspired by traditional topic models. For data exploration and inductive coding, LLMs can surface novel patterns and assist in the initial exploration of themes and codes, potentially enabling a deeper understanding. However, while LLMs can replicate many primary themes identified by humans, they often miss more subtle and complex interpretations (Morgan, 2023; Zhang et al., 2023; Torii et al., 2024). Hamilton et al. (2023) observed complementary patterns between human and ChatGPT-generated codes, with each identifying themes that the other did not. Barany et al. (2024) investigated the use of ChatGPT for codebook creation, finding hybrid approaches to be more reliable. They suggest LLMs can contribute to the improvement of codebooks, but stress that human involvement remains crucial. While combining interpretive and positivist methods can cause epistemic tensions and discussions among researchers (Baumer et al., 2017; Muller et al., 2016; Paulus and Marone, 2024), studies highlight mutual compatibility (Baumer et al., 2017; Chatzichristos, 2025; Gauthier and Wallace, 2022; Lam et al., 2024; Muller et al., 2016). AI also holds potential for mediating discussions and facilitating collaboration by highlighting coding discrepancies, and suggesting resolutions (Gao et al., 2023; Ganji et al., 2018; Drouhard et al., 2017).
@@
While these efforts illustrate potential, they also expose tensions. Concerns persist around bias (Bail, 2024; Törnberg, 2023b; Ziems et al., 2024; Anis and French, 2023; Owoahene Acheampong and Nyaaba, 2024; Bano et al., 2023), lack of reliability for non-English languages (Heseltine and Clemm von Hohenberg, 2024; Suter and Meckel, 2024), and the risk of treating AI as “culturally agnostic” despite its privileging of certain perspectives (Paulus and Marone, 2024). Such issues are particularly important when engaging minority groups or Global South contexts, where AI risks reproducing inequalities in representation and missing interpretative nuance (Kumar and Karusala, 2021; Wacharamanotham et al., 2020; Rifat et al., 2023; Jang et al., 2023; Okolo et al., 2022). This raises concerns regarding bias in interpretation, but also research ethics of fairness and inclusion (Kempny et al., 2025). Moreover, perspectives from the Global Souths are largely absent from AI development, and little is known on how AI is used and deployed in the Global Souths (Okolo et al., 2022; Linxen et al., 2021; Tahaei et al., 2024).
Another tension arises from the very foundations of AI itself: borne on statistical methods and probability, AI reflects a positivist paradigm that can impede the interpretative nature of qualitative inquiry (Chatzichristos, 2025; Paulus and Marone, 2024; Jiang et al., 2021). This is amplified by the problem of overreliance on AI, facilitated by the plausibility of AI generations: in QDA, AI-generated ”interpretations” or suggestions may be simply accepted, instead of being critically interrogated (Kempny et al., 2025; Okolo et al., 2022; Buçinca et al., 2021), creating lopsided structures of collaboration between AI and human researchers in QDA. This poses specific challenges for teams involving collaborators with varying degrees of literacy regarding AI as well as QDA, especially in interdisciplinary fields like CSCW and HCI. Thus, while AI holds potential to bridge “big” and “small” data approaches and to enhance efficiency, it raises equally urgent questions about reliability, ethics, and researcher agency. These risks and tensions require careful considerations for designing human-AI collaboration in QDA, preserving the depth and richness of QDA while leveraging the potentials of AI.
2.4. Designing for Human-AI Collaboration in QDA
@@
Integrating AI into QDA is not simply a matter of automation, but of collaboration as well: It requires careful design for human–AI and human-human collaboration in interpretive, context-sensitive work (Jiang et al., 2021). Existing tools vary widely in how they support AI involvement, from minimal assistance to active roles in coding and analysis, with different levels of human involvement. Moreover, designing for researchers differing in proficiency in QDA, tool usage, as well as AI capability (Lubars and Tan, 2019; Feuston and Brubaker, 2021) poses specific challenges, especially in diverse fields such as CSCW and HCI. For example, Gauthier and Wallace (2022) developed a toolkit to facilitate computational thematic analysis, emphasizing the need to design it accessibly for a community with different proficiencies in tool usage as well as AI. Thus, they call for ethical designs to involve friction to remind researchers of the imperfection of computational models and issues concerning ethics and transparency. These challenges are also identified by Kapania et al. (2025) who found HCI researchers struggling to identify such issues in their projects and needing guidance. Similarly, Lam et al. (2024) suggest tools to be transparent to the researcher and acknowledging their limitations, e. g., by showcasing explicit inclusion criteria and scores. At the same time, designs should avoid creating excessive “unanticipated labor” for researchers (Feuston and Brubaker, 2021).
@@
While prior studies provide valuable insights into the potentials of AI in QDA (Tai et al., 2024; Morgan, 2023; Rasheed et al., 2024; Lam et al., 2024; Gauthier and Wallace, 2022), there is still limited understanding of how AI can be responsibly integrated into researchers’ real-life QDA workflows, respecting their concerns and preferences (Lubars and Tan, 2019; Feuston and Brubaker, 2021), facilitating collaboration between researchers as well as researchers and AI, while consolidating different research epistemologies. We are thus in need for comprehensive frameworks addressing these issues (Owoahene Acheampong and Nyaaba, 2024; Kapania et al., 2025). Moreover, it remains crucial to critically reflect not only on AI capability, but also what it should do in QDA, preserving researcher concerns and preferences (Feuston and Brubaker, 2021; Lubars and Tan, 2019). We extend this to the era of LLMs, identifying new steps, tasks, and different levels of involvement of AI to support QDA while maintaining methodological integrity. To address these gaps, we study HCI researchers’ QDA workflows and their perspectives on the involvement of AI in qualitative research. Specifically, we investigate the following research questions:
@@ 3.2. Recruitment & Participants
Table 1 summarizes the participant sample. We interviewed 15 HCI researchers with backgrounds in computer science, cybersecurity, engineering, and social sciences, all firmly embedded in HCI research, with some specifying to focus on USP. Twelve participants were early career researchers (mostly PhD candidates), and three were senior researchers. Eight had worked on 1-3 QDA projects, four on 4-6, and three on more than 7 projects, with QDA experience ranging from 1 to 9 years. We recruited participants until data saturation was reached. We reached data saturation around interview 12 and conducted three additional interviews to confirm this.
@@ 4.1.2. Uncovering Participants’ Actual Coding Workflow
Figure 1. Detailed overview of participants’ real-world coding workflow, consisting of three stages for codebook creation, codebook refinement, codebook application, and different steps taken in inductive and deductive approaches.
The figure shows a flowchart depicting the 3 stages of the coding workflow: Codebook Creation, Codebook Refinement, and Codebook Application. It has 2 starting points: 1 for inductive, and 2 for deductive, and 1 endpoint labeled Done. 12 elements depict different steps in this process, including visuals representing individual or joint coding. 2 types of lines connect the different elements: arrows depict the element as next step. Dotted arrows depict the next element as optional. The elements representing steps are listed below. Starting point 1: Inductive Step 1: Cluster/Inspect Data. Individual or joint. From this, an arrow leads through pictograms of different data sets to step 2. Step 2: An element with multiple pictograms labeld Individual and 1 labeled Joint are depicted. They are connected to elements labeled Codebook 1 to Codebook N. An arrow leads to the step 3. Step 3. An element labeled First Full Codebook is shown at the end of the stage Codebook Creation. An arrow leads to 3 different elements: Step 3.1, Step 3.2, and Step $3.$ A dotted arrow leads to this arrow, labeled Optional IRR. Step 3.1 (Re-)Apply Codebook in the Stage of Codebook Application. An arrow leads to the next element: Refine Codebook, including labels for Remove Codes and Merge Codes. A dotted arrow labeled IRR high leads to the final step: Done. From this dotted arrow, another arrow with the description IRR low leads back to the element Refine Codebook. Another arrow labeled never ending process leads back to the element (Re-)Apply Codebook. Step 3.2: Test Codebook in the Stage of Codebook Refinement. This element consists of 3 pictograms for coders. 1: Individual, 2: Joint, 3: Joint with new. An arrow leads to step 4. Step 4: Refine Codebook in the Stage of Codebook Refinement. The element contains labels for Remove codes and Reduce granularity. An arrow leads to step 5. Step 5. (Re-)Apply Codebook in the stage of Codebook Application. It contains labels for Rest of data and Seen data. It further shows iterations labeled Single, Multiple, or Random Multiple. A dotted arrow labeled IRR high leads back to the same element (Re-)Apply Codebook. A dotted arrow labeled IRR low leads back to step 4: Refine Codebook. Another dotted arrow leads to the last element: Done. Step 3.3: An element labeled Discuss with Coded Data. From here, an arrow leads to step 5: (Re-)Apply Codebook. Starting point 2: Deductive Step 1: An element labeled Deduce from Literature / Research Questions. A dotted arrow leads to Step 2 in Inductive Coding. A straight arrow leads to the next step in Codebook Application. Step 2.2: An element labeled Apply Codebook. A dotted arrow leads to the last element: Done.